OpenStudy (anonymous):

At a restaurant, Brandy paid twice as much as Katie, and Katie paid $3 more than Stephanie. If the total combined amount they paid was $56, how much did Katie pay?

OpenStudy (dangerousjesse):

So \(x+3+(x\times 2)+x=56\)

OpenStudy (dangerousjesse):

\(4x+3=56\)

OpenStudy (dangerousjesse):

Subtract three.. \(4x=53\)

OpenStudy (dangerousjesse):

Divide by four then add three to get your answer